# Input Convex Lipschitz RNN: A Fast and Robust Approach for Engineering Tasks

### v5

@@ 5.2 Modeling Performance
5.3 MPC Performance
Next, incorporating an ICNN (e.g., ICLRNN) into the aforementioned MPC in Eq. (2) transforms it into a convex optimization problem. Subsequently, PyIpopt, as a Python version of IPOPT (Wächter and Biegler, 2006), is used to address the MPC problem. The integration time step is set at $h_{c}=1\times 10^{-4}~{}hr$, while the sampling period remains $\Delta=5\times 10^{-3}~{}hr$. Additionally, we introduce a Lyapunov constraint to the system, as follows:
$V({\widetilde{\bm{\mathbf{x}}}}(t))<V(\bm{\mathbf{x}}(t_{k})),\forall t\in[t_{k},t_{k+N})$ (22)
where the Lyapunov-based constraint $V$ is used to ensure closed-loop stability within MPC by mandating a decrease in the value of $V(x)$ [^4] Lyapunov functions $V(x)$ are analogous to energy functions in some sense that the system is stable at the lowest energy point where $V(x)=0$ at $x=0$.. The design of the Lyapunov function, denoted as $V(\bm{\mathbf{x}})=\bm{\mathbf{x}}^{\top}\bm{\mathbf{P}}\bm{\mathbf{x}}$, involves configuring the positive definite matrix $\bm{\mathbf{P}}$ as $\begin{bmatrix}1060&22\\ 22&0.52\end{bmatrix}$, which ensures the convexity of the MPC.
We execute MPC for a fixed timeframe on an Intel Core i7-12700 processor with 64 GB of RAM for various models and repeat with various initial conditions to evaluate the performance of different NN-based MPC. Fig. 3 shows one of the initial conditions, from which it is evident that the ICLRNN-based MPC converges, whereas other NN-based MPC models do not. The fastest convergence speed highlights the superior computational efficiency of the ICLRNN-based MPC.
Figure 3: MPC converging path in a fixed timeframe with an initial condition at $[-1.5~{}kmol/m^{3},70~{}K]$.
Moreover, it is important to highlight that ICLRNN-based MPC consistently demonstrates the fastest convergence speed across various initial conditions, thereby surpassing the state-of-the-art methodologies. Figs. 4 and 5 provide compelling evidence that the ICLRNN-based MPC achieves convergence, unlike other NN-based MPC, within a fixed timeframe. This outcome distinctly indicates that ICLRNN-based MPC exhibits the fastest convergence speed under identical settings.
In Fig. 4, the left panels show the results corresponding to the initial condition set at [1.5[-1.5~{}[ - 1.5 $kmol/m^{3},70~{}K]$. Specifically, the top left panel displays the concentration profile, the middle left panel shows the temperature, and the bottom left panel exhibits the converging path. Additionally, the right panels showcase the results from the initial condition at $[1.5~{}kmol/m^{3},-70~{}K]$, featuring the concentration profile in the top right panel, temperature profile in the middle right panel, and the converging path in the bottom right panel. Moving to Fig. 5, the left panels depict the results obtained from the initial condition of $[-1.25~{}kmol/m^{3},50~{}K]$. Here, the top left panel represents the concentration, the middle left panel illustrates the temperature, and the bottom left panel shows the converging path. Similarly, the right panels exhibit the results of the initial condition at $[1.25~{}kmol/m^{3},-50~{}K]$. In summary, across various initial conditions, ICLRNN-based MPC consistently demonstrates the fastest convergence speed, thereby surpassing the state-of-the-art methods.
Figure 4: Concentration profile (top left), temperature profile (middle left), converging path (bottom left) in a fixed timeframe with an initial condition at $[-1.5~{}kmol/m^{3},70~{}K]$ under MPC, and concentration profile (top right), temperature profile (middle right), converging path (bottom right) in a fixed timeframe with an initial condition at $[1.5~{}kmol/m^{3},-70~{}K]$.
Figure 5: Concentration profile (top left), temperature profile (middle left), converging path (bottom left) in a fixed timeframe with an initial condition at $[-1.25~{}kmol/m^{3},50~{}K]$, and concentration profile (top right), temperature profile (middle right), converging path (bottom right) in a fixed timeframe with an initial condition at $[1.25~{}kmol/m^{3},-50~{}K]$.
6 Application to Real-World Solar Irradiance Prediction
In this section, we aim to forecast solar irradiance using real-time data for LHT Holdings, a wood pallet manufacturing industry based in Singapore.
6.1 System Description and Modeling
Before the installation of the solar photovoltaic (PV) system, LHT Holdings relied solely on the main utility grid to fulfill its energy needs. The solar PV system was successfully installed by 10 Degree Solar in late 2022 (the manufacturing pipeline of LHT Holdings and the solar PV system is described in our previous work (Wang et al., 2025)). This transition to solar energy was primarily driven by the aim of minimizing manufacturing costs. Three significant uncertainties were identified within the industry, namely solar irradiance affecting the solar PV system’s efficiency, the dynamic real-time pricing of energy sourced from the main utility grid, and the fluctuating energy demand stemming from unexpected customer orders.
Currently, the industrial facility relies primarily on solar PV as its main energy source, with the main utility grid serving as a backup. Excess solar energy is stored for later use. Accurate solar irradiance forecasting is crucial for efficient energy planning and cost reduction. Using sensors installed by the Solar Energy Research Institute of Singapore (SERIS) and Singapore Institute of Manufacturing Technology (SIMTech), including irradiance, humidity, wind, and temperature sensors, data is systematically recorded on a minute-by-minute basis and seamlessly uploaded online into the system. Alongside data similar to that used by Zelikman et al. (2020), we predict minute-by-minute solar irradiance to optimize energy dispatch. In this time-series forecasting task, all models undergo training and evaluation using identical settings. These settings encompass model size of (256, 256), MSE as the loss function, and the Adam Optimizer. For the training and validation phases, we use data spanning from January 14, 2023, to December 16, 2023. Subsequently, data from December 17, 2023, to January 1, 2024, is exclusively reserved for testing and evaluating the models’ predictive performance.
Furthermore, from an engineering perspective, the training and implementation of neural network models are often constrained by limited computational resources, which are typically not as ample as those in computer science. Consequently, engineering applications require relatively simple models capable of effectively handling challenges such as data noise and convexity. To ensure the selection of the most suitable model, it is essential to conduct comparisons under consistent settings (i.e., the same model complexity). This approach ensures fairness, as we evaluate not only accuracy, but also other critical factors such as model complexity, which can be quantified by FLOPs.
6.2 Performance Analysis for Solar Irradiance Forecasting
We leverage FLOPs as a measure to assess the model complexity. As depicted in Table 4, it is evident that among all the models, ICLRNN demonstrates the most efficient computational performance as a result of its lowest FLOPs (excluding plain RNN). Consequently, this superior efficiency positions ICLRNN as the top-performing model when considering computational resources.
Table 4: Hyperparameters of neural network models for solar irradiance forecasting
@@ 6.2 Performance Analysis for Solar Irradiance Forecasting
Plain RNN Tanh $\bm{\mathbf{x}}^{\top}_{t}$ 399,362
Plain LSTM Tanh $\bm{\mathbf{x}}^{\top}_{t}$ 1,596,418
LRNN Tanh $\bm{\mathbf{x}}^{\top}_{t}$ 1,195,010
ICRNN ReLU $\hat{\bm{\mathbf{x}}}_{t}=\begin{bmatrix}\bm{\mathbf{x}}^{\top}_{t},-\bm{\mathbf{x}}^{\top}_{t}\end{bmatrix}^{\top}$ 2,498,562
ICLRNN (Ours) ReLU $\hat{\bm{\mathbf{x}}}_{t}=\begin{bmatrix}\bm{\mathbf{x}}^{\top}_{t},-\bm{\mathbf{x}}^{\top}_{t}\end{bmatrix}^{\top}$ 401,410
Given the real-world context of this task, introducing artificial noise for evaluating non-adversarial robustness becomes unnecessary as the inherent noise within real-world data suffices. Notably, our observations reveal that ICRNN encounters challenges with training when the hypothesis space size is large (i.e., the number of hidden neurons per layer is more than 128), thereby hindering its ability to effectively model complex system dynamics. While RNN and LSTM exhibit a limited capacity to capture system dynamics, they struggle notably when faced with sudden changes in solar irradiance, leading to less accurate predictions.
In Fig. 6(a), we present results from the best state-of-the-art model, LRNN, alongside our model’s outcomes specifically for December 28, 2023. The LRNN, constrained by Lipschitz continuity, exhibits significantly improved performance compared to the RNN and LSTM models. However, it still encounters challenges in accurately predicting solar irradiance during sudden fluctuations, and its excessively high FLOPs signify its computational intensity (i.e., 2,498,562 for LRNN compared to 401,410 for ICLRNN). Moreover, slight fluctuations in the LRNN predictions occur during periods of minimal solar irradiance at the start and end of the day. Highlighted within Fig. 6(b), our ICLRNN outperforms LRNN and other state-of-the-art models, particularly in accurately predicting sudden changes in solar irradiance, which highlights the superior predictive capability of ICLRNN, showcasing its effectiveness in capturing and predicting nuanced fluctuations within the solar irradiance data. Notably, these observations persist across all trials and are not specific solely to December 28, 2023.
(a) LRNN performances on 2023-12-28 for solar irradiance prediction.
(b) ICLRNN performances on 2023-12-28 for solar irradiance prediction.
Figure 6: Solar irradiance forecasting performance of LRNN and ICLRNN on 2023-12-28.
Specifically, Fig. 7 shows the model performance on solar irradiance prediction on December 28, 2023. The top two figures display the RNN and LSTM prediction outcomes, respectively. Both models demonstrate poor performance, failing to adjust adequately when faced with sudden changes in solar irradiance, leading to overshooting. Moreover, the middle two figures reveal that the ICRNN encounters issues related to the exploding gradient problem, particularly when the hypothesis space size becomes excessively large. The middle left figure shows the prediction performance of ICRNN with 128 hidden neurons per layer, while the middle right figure shows the prediction performance of ICRNN with 256 hidden neurons per layer. When surpassing 128 hidden neurons per layer, the ICRNN struggles to effectively capture and learn the system’s dynamics. Notably, these observations persist across all trials and are not specific solely to December 28, 2023. It is clearly shown in Fig. 7 that our ICLRNN model surpasses state-of-the-art models, demonstrating superior predictive capabilities across various scenarios.
Figure 7: RNN (top left), LSTM (top right), ICRNN (middle left and middle right with neural architecture of (128, 128) and (256, 256) respectively), LRNN (bottom left), and ICLRNN (bottom right) performance on 2023-12-28 for solar irradiance prediction.

### v6

@@ 5.2 Modeling Performance
5.3 Computational Efficiency in MPC for CSTR Control
In addition to robustness against noise, model computational efficiency is also a significant factor in process modeling, optimization, and control applications. In this work, we employ the framework of MPC to evaluate the models’ computational efficiency, which is an advanced control methodology widely implemented in industrial chemical plants and energy systems. Specifically, a typical MPC consists of model prediction in the horizons of a fixed temporal window, optimal solving of a properly set objective function, and actuator correction by updating the first horizon’s optimal solution. Therefore, the computational efficiency of MPC largely depends on how rapidly the optimal solution can be obtained, which presents challenges to the model’s tractability in optimization problems, especially those of systems with strong nonlinearity.
[…]
Figure 3: Closed-loop state profiles under neural network model-based MPC.
Table 4: MPC computation time of neural network models
Model Average computation time (s)
Plain RNN 152.254
Plain LSTM 149.494
LRNN 141.467
ICRNN 81.779
ICL-RNN (Ours) 100.997
The average computation time of ICLRNN-MPC is 33.67%, 32.44%, and 28.61% lower than that of RNN-MPC, LSTM-MPC, and LRNN-MPC, respectively, and is 19.03% higher than that of ICRNN-MPC. This advantage in the computational efficiency of ICLRNN and ICRNN is due to the convex optimization introduced in their training process.
Furthermore, while the introduction of the Lipschitz constraint makes the computational efficiency of ICLRNN-MPC slightly lower than that of ICRNN-MPC, it effectively enhances the robustness against noise (as shown in Sec. 5.2) without significantly compromising the computational efficiency. In this case, the proposed ICLRNN integrates strong robustness with high computational efficiency by integrating convex optimization and Lipschitz constraint. This synergy provides practical value in real-world process control, where models’ tolerance for data noise and computational efficiency during operation are both highly required.
6 Modeling and Control of ORC-Based Waste Heat Recovery System
6.1 ORC-Based Waste Heat Recovery System
In this section, we consider an ORC-based waste heat recovery system, a more complex example in energy systems, to show the performance of the proposed ICLRNN. Specifically, the ORC system is designed to recover the energy of waste heat and convert it into mechanical power. A typical ORC system mainly comprises an evaporator, pump, reservoir, condenser, and expander, through which the organic working fluid R245fa circulates to carry out the energy conversion process. The dynamics of the system can be summarized as follows: The pump transfers the liquid R245fa to the evaporator, where it absorbs heat from the waste heat source and undergoes phase transition into vapor, which then expands through the expander to generate the shaft work. Subsequently, the vapor is condensed back to liquid form in the air-cooled condenser and recirculated to the evaporator, thereby completing the ORC circulation.
The first-principles model of the ORC system adopted in this study is based on (Shi et al., 2023), The reservoir outlet temperature and pressure are considered to be constants due to the assumption of a sufficiently large reservoir. Furthermore, given that the condenser is typically subject to negligible disturbances in practical operation, the overall system dynamics is predominantly governed by the heat and mass transfer processes within the pump, expander, and most critically, the evaporator. In this case, the models of the expander and pump are given as follows:
$\dot{m}_{exp}$ $=\eta_{v,exp}\rho_{o,ex}V_{exp}\omega_{exp}$ (22a)
$h_{o,exp}$ $=h_{i,exp}+\eta_{ise,exp}(h_{ise,exp}-h_{i,exp})$ (22b)
$\dot{m}_{pump}$ $=\eta_{v,pump}\rho_{o,pump}V_{pump}\omega_{pump}$ (23a)
$h_{o,pump}$ $=h_{i,pump}+\frac{1}{\eta_{ise,pump}}(h_{ise,pump}-h_{i,pump})$ (23b)
where $\dot{m}$ denotes the mass flow rates through the expander and pump, with subscripts “exp” and “pump” representing the expander and pump, respectively. $\eta_{v}$, $\rho_{o}$, $V$, and $\omega$ are the corresponding volumetric efficiency, outlet fluid density, displacement volume, and rotational speed, respectively; $h_{i}$ and $h_{o}$ represent the inlet and outlet specific enthalpy, while $h_{ise}$ and $\eta_{ise}$ denote the isentropic outlet enthalpy and efficiency. Regarding crucial evaporator dynamics, the moving boundary (MB) modeling method used in (Shi et al., 2023) divides the evaporator into three regions (subcooled, two-phase, superheated), corresponding to the phase states of the working fluid along the evaporator. Consequently, the mass and energy conservation equations for each region ($k=1,2,3$ represent the subcooled, two-phase, superheated regions of the evaporator, respectively) are formulated as follows:
$\frac{\partial A_{eva}\rho^{k}_{eva}}{\partial t}=\frac{\partial\dot{m}^{k}_{eva}}{\partial z}$ (24)
$\frac{\partial\dot{m}^{k}_{eva}h^{k}_{eva}}{\partial z}=-\frac{\partial(A_{eva}\rho^{k}_{eva}h^{k}_{eva}-A_{eva}\rho^{k}_{eva})}{\partial t}+\pi\alpha^{k}_{i,eva}D^{k}_{i,eva}(T^{k}_{w,eva}-\bar{T}^{k}_{r,eva})$ (25)
where $A_{eva}$ is the flow cross-sectional area of the evaporator channel. $\rho^{k}_{eva}$, $\dot{m}^{k}_{eva}$, and $h^{k}_{eva}$ are the fluid density, mass flow rate, and enthalpy in region $k$, respectively. $D^{k}_{i,eva}$ denote the inner hydraulic diameter in region $k$, $T^{k}_{w,eva}$ and $\bar{T}^{k}_{r,eva}$ represent the wall temperature and mean fluid temperature in region $k$, while $z$ and $t$ are defined as the length coordinate and time. $\alpha^{k}_{i,eva}$ is the inner heat transfer coefficient in region $k$, which is obtained based on the local mass flow rates and phase distribution in each region. Specifically, the coefficients for the sub-cooled ($k=1$) and superheated ($k=3$) regions are scaled from reference values:
$\alpha^{1}_{i,\mathrm{eva}}$ $=\alpha^{1}_{i0,eva}\left(\frac{\dot{m}_{i,\mathrm{eva}}}{\dot{m}_{f_{0}}}\right)^{0.8}$ (26a)
$\alpha^{3}_{i,\mathrm{eva}}$ $=\alpha^{3}_{i0,eva}\left(\frac{\dot{m}_{o,\mathrm{eva}}}{\dot{m}_{f_{0}}}\right)^{0.9}$ (26b)
where $\dot{m}_{i,eva}$ and $\dot{m}_{o,\mathrm{eva}}$ are the liquid and vapor mass flow rates, respectively; $\alpha^{1}_{i0,\mathrm{eva}}$ and $\alpha^{3}_{i0,\mathrm{eva}}$ denote the reference inner heat transfer coefficients in the sub-cooled and superheated regions, while $\dot{m}_{f0}$ is the reference mass flow rate. Meanwhile, $\alpha^{2}_{i,\mathrm{eva}}$ for the two-phase region is calculated as a semi-empirical function of the local phase fraction $x_{e}$, the inner heat transfer coefficients of the neighboring regions, and the gas-to-liquid density ratio $d=\rho_{g}/\rho_{l}$.
$\alpha^{2}_{i,\mathrm{eva}}=\alpha^{1}_{i,\mathrm{eva}}\Bigg[\big((1-x_{e})+1.2x_{e}^{0.4}(1-x_{e})d^{0.37}\big)^{-2.2}+\bigg(\frac{\alpha^{3}_{i,\mathrm{eva}}}{\alpha^{1}_{i,\mathrm{eva}}}x_{e}^{0.01}(1+8(1-x_{e})^{0.7})d^{0.67}\bigg)^{-2}\Bigg]^{-0.5}$ (27)
In addition, the energy conservation of the pipe wall is also introduced to help increase the model accuracy:
$(C_{p}\rho A)_{w,eva}\frac{\partial T^{k}_{w,eva}}{\partial t}=\pi(D^{k}_{o,eva}\alpha_{o,eva}(T^{k}_{a,eva}-T^{k}_{w,eva})+D^{k}_{i,eva}\alpha^{k}_{i,eva}(\bar{T}^{k}_{r,eva}-T^{k}_{w,eva}))$ (28)
where $(C_{p}\rho A)_{w,eva}$ represents the thermal capacitance of the evaporator pipe wall, defined as the product of the wall’s specific heat at constant pressure $C_{p}$, density $\rho$, and cross-sectional area $A$. $D^{k}_{o,eva}$ and $T^{k}_{a,eva}$ are the outer diameter of the evaporator pipe and heat source temperature in region $k$, respectively, while $\alpha_{o,eva}$ is the outer heat transfer coefficient, obtained by scaling the reference outer coefficient according to the exhaust air mass flow rate:
$\alpha_{o,\mathrm{eva}}=\alpha_{o0}\left(\frac{\dot{m}_{a}}{\dot{m}_{a_{0}}}\right)^{0.5}$ (29)
where $\alpha_{o0}$ and $\dot{m}_{a0}$ are the reference outer heat transfer coefficient and reference exhaust air mass flow rate, respectively. Then, the heat source outlet temperature in region $k$ is presented as:
$T^{k}_{o,eva}=e^{-\phi}T^{k}_{i,eva}+(1-e^{-\phi})T^{k}_{w,eva}$ (30)
$\phi=\frac{A^{k}_{tr,eva}\alpha_{o,eva}}{\dot{m}_{a}C_{a}}$ (31)
where $T^{k}_{o,eva}$ and $T^{k}_{i,\mathrm{eva}}$ denote the outlet and inlet temperature in region $k$, respectively. $A^{k}_{\mathrm{tr,eva}}$ is the heat transfer area between the waste heat and the pipe wall in region $k$, and $\dot{m}_{a}$ and $C_{a}$ are the mass flow rate and specific heat capacity of the heat source, respectively. Following that, the lengths of the sub-cooled and two-phase regions, $L^{1}_{eva}$ and $L^{2}_{eva}$, as well as the evaporator pressure $P_{eva}$ and superheat $SH$, are calculated by integrating Eqs. (24) and (25) with the Leibniz formula, as expressed in Eq. (32):
$\int_{z_{1}}^{z_{2}}\frac{\partial f(z,t)}{\partial t}\,dz=\frac{d}{dt}\int_{z_{1}}^{z_{2}}{\partial f(z,t)}\,dz+f(z_{1},t)\frac{dz_{1}}{dt}-f(z_{2},t)\frac{dz_{2}}{dt}$ (32)
and obtaining the intermediate physical properties (including evaporation pressure at the two-phase region and the saturation temperature at the evaporator outlet) through Refprop 9.1, thereby capturing the system dynamics. On the basis of the above-mentioned relations, the first-principles model of the ORC system is constructed as a seven-state ODE:
$D_{eva}\dot{x}_{eva}=f_{eva}(x_{eva},u_{eva})$ (33)
where $x_{eva}$ denotes the selected state variables $L^{1}_{eva}$, $L^{2}_{eva}$ $T^{1}_{w,eva}$, $T^{2}_{w,eva}$, $T^{3}_{w,eva}$, $P_{eva}$, $SH$], while $u_{eva}$ denotes the input control actions $\omega_{exp}$, $\omega_{pump}$]. Therefore, the inputs of the ORC system in this study consist of $\omega_{exp,t}$, $\omega_{pump,t}$, $L^{1}_{eva,t}$, $L^{2}_{eva,t}$, $T^{1}_{w,eva,t}$, $T^{2}_{w,eva,t}$, $T^{3}_{w,eva,t}$, $P_{eva,t}$, $SH_{t}$] at the current time step $t$, while the outputs $L^{1}_{eva,t+1}$, $L^{2}_{eva,t+1}$ $T^{1}_{w,eva,t+1}$, $T^{2}_{w,eva,t+1}$, $T^{3}_{w,eva,t+1}$, $P_{eva,t+1}$, $SH_{t+1}$] at time step $t+1$ entail the state trajectory over the subsequent $n$ time steps, which is set to $5$ in this case.
Table 5: Parameter values for the ORC model
Parameter Value Parameter Value
$V_{\mathrm{exp}}$ $1.4\times 10^{-4}~\mathrm{m^{3}}$ $\eta_{v,\mathrm{exp}}$ $0.6$
$V_{\mathrm{pump}}$ $1.3\times 10^{-5}~\mathrm{m^{3}}$ $\eta_{v,\mathrm{pump}}$ $0.6$
$A_{\mathrm{eva}}$ $3.142\times 10^{-4}~\mathrm{m^{2}}$ $D_{i,\mathrm{eva}}$ $0.02~\mathrm{m}$
$D_{o,\mathrm{eva}}$ $0.022~\mathrm{m}$ $A_{w,\mathrm{eva}}$ $6.91\times 10^{-5}~\mathrm{m^{2}}$
$\alpha_{o0,\mathrm{eva}}$ $435~\mathrm{W/(m^{2}\cdot K)}$ $\alpha^{1}_{i0,\mathrm{eva}}$ $1090.0~\mathrm{W/(m^{2}\cdot K)}$
$\alpha^{3}_{i0,\mathrm{eva}}$ $661.2~\mathrm{W/(m^{2}\cdot K)}$ $\dot{m}_{f_{0}}$ $0.132~\mathrm{kg/s}$
$\dot{m}_{a_{0}}$ $0.200~\mathrm{kg/s}$ $C_{p,w,\mathrm{eva}}$ $0.385~\mathrm{kJ/(kg\cdot K)}$
$\rho_{w,\mathrm{eva}}$ $8.96\times 10^{3}~\mathrm{kg/m^{3}}$ $L^{1}_{eva}+L^{2}_{eva}+L^{3}_{eva}$ $9~\mathrm{m}$
Similarly, open-loop simulations of this ORC system are also performed using the explicit Euler method to generate data that mimics the real-world data for neural network training. The detailed system parameters are summarized in Table 5.
6.2 System Modeling
To further evaluate the model’s robustness against noise in the application of the ORC system, Gaussian noise is also added to the training data (i.e., output trajectory $\bm{\mathbf{y}}$) in an additive manner following Eq. (21). Meanwhile, all models are trained and evaluated using consistent configurations, including MinMax scaling (scaling the data to a range between 0 and 1), a batch size of 256, a two-layer neural architecture with 64 neurons per layer (i.e., model size of (64, 64)), the mean-squared error (MSE) loss function, and the Adam optimizer. Detailed hyperparameter values, including the FLOPs for quantifying model complexity, are provided in Table 6. Based on FLOPs, the model complexity ranks as follows: RNN << ICL-RNN << ICRNN << LRNN << LSTM. Furthermore, it is important to note that all experiments were conducted three times. In the plots, the solid line represents the mean across the three trials on different unseen reactions, while the shaded region illustrates the standard deviation.
Table 6: Hyperparameters of neural network models
@@ 6.2 System Modeling
Plain RNN Tanh $\bm{\mathbf{x}}^{\top}_{t}$ 30, 499
Plain LSTM Tanh $\bm{\mathbf{x}}^{\top}_{t}$ 260, 736
LRNN Tanh $\bm{\mathbf{x}}^{\top}_{t}$ 162, 083
ICRNN ReLU $\hat{\bm{\mathbf{x}}}_{t}=\begin{bmatrix}\bm{\mathbf{x}}^{\top}_{t},-\bm{\mathbf{x}}^{\top}_{t}\end{bmatrix}^{\top}$ 82, 979
ICL-RNN (ours) ReLU $\hat{\bm{\mathbf{x}}}_{t}=\begin{bmatrix}\bm{\mathbf{x}}^{\top}_{t},-\bm{\mathbf{x}}^{\top}_{t}\end{bmatrix}^{\top}$ 30, 755
The modeling results are comparable to those observed in the CSTR case, including the analysis and conclusions. From Fig. 4, for the testing MSE, the $y$-axis of the left figure is presented in a linear scale, while the right figure uses a logarithmic scale. All three models - ICRNN, LRNN, and ICL-RNN - demonstrate robustness to noise. However, in terms of modeling accuracy, the performance ranks as LRNN >> ICL-RNN >> ICRNN. Nonetheless, the advantage of the input convex property becomes evident in controlling neural network-based MPC.
(a) Linear scale
(b) Logarithmic scale
Figure 4: Testing MSE vs. degree of additive noise over 3 random trials.
(a) Linear scale
(b) Logarithmic scale
Figure 5: Lipschitz constant of the trained model vs. degree of additive noise over 3 random trials.
Similarly, from Fig. 5, for the Lipschitz constant, the $y$-axis of the left figure is presented in a linear scale, while the right figure uses a logarithmic scale. Due to the recurrent nature of the neural networks, obtaining a precise Lipschitz constant is challenging. Therefore, a numerical method was used to approximate it (see Remark 2). For all three models - ICRNN, LRNN, and ICL-RNN - the Lipschitz constant is less than or equal to 1, with values that are approximately the same, while any discrepancies observed are due to approximation errors.
The ICRNN model becomes untrainable when the number of neurons per layer exceeds 256. In contrast, our ICL-RNN remains trainable, as the Lipschitz constraint stabilizes the training process (see Table 7). This stabilization property, achieved through spectral normalization, has been previously observed and utilized in stabilizing the training of GANs in generative AI (Miyato et al., 2018), but it has not yet been widely discussed in the context of RNNs.
Table 7: Testing MSE of ICRNN and ICL-RNN with respect to increasing hypothesis space
Testing MSE Lipschitz Constant
Model Size ICRNN ICL-RNN ICRNN ICL-RNN
(32, 32) $8.913\times 10^{-3}\pm 1.14\times 10^{-4}$ $8.697\times 10^{-3}\pm 8.81\times 10^{-5}$ $5.03\times 10^{-1}\pm 1.03\times 10^{-1}$ $5.03\times 10^{-1}\pm 8.50\times 10^{-2}$
(64, 64) $8.926\times 10^{-3}\pm 7.83\times 10^{-4}$ $8.636\times 10^{-3}\pm 9.13\times 10^{-5}$ $6.62\times 10^{-1}\pm 1.85\times 10^{-1}$ $5.11\times 10^{-1}\pm 8.41\times 10^{-2}$
(128, 128) $1.944\times 10^{-2}\pm 2.69\times 10^{-4}$ $8.612\times 10^{-3}\pm 9.22\times 10^{-5}$ $7.92\times 10^{-1}\pm 5.03\times 10^{-2}$ $5.30\times 10^{-1}\pm 8.59\times 10^{-2}$
(256, 256) $2.304\pm 2.29\times 10^{-2}$ $8.904\times 10^{-3}\pm 1.13\times 10^{-5}$ $3.352\pm 1.262$ $4.07\times 10^{-1}\pm 5.40\times 10^{-2}$
(512, 512) NaN $8.046\times 10^{-3}\pm 7.79\times 10^{-6}$ NaN $4.90\times 10^{-1}\pm 1.60\times 10^{-2}$
6.3 Computational Efficiency in MPC for ORC Control
In this section, we evaluate the computational efficiency of MPC using different models for the ORC system. This case study represents a more complex and practically representative benchmark, allowing us to further validate the computational efficiency enhancement achieved by the proposed ICLRNN and examine whether these advantages generalize to systems with different dynamics characteristics. Specifically, the initial states of the ORC system are shown in Table 8, among which $SH$ and $P_{eva}$ are the control objectives with the setpoints, $SH_{S}$ and $P_{eva,s}$, set to 14.83 K and 2287.03 kPa, respectively.
Table 8: Initial States of the ORC system
Initial State Value Initial State Value
$T^{1}_{w,\mathrm{eva}}$ $424.78~\mathrm{K}$ $T^{2}_{w,\mathrm{eva}}$ $433.23~\mathrm{K}$
$T^{3}_{w,\mathrm{eva}}$ $474.75~\mathrm{K}$ $L^{1}_{\mathrm{eva}}$ $4.14~\mathrm{m}$
$L^{2}_{\mathrm{eva}}$ $3.47~\mathrm{m}$ $P_{\mathrm{eva}}$ $2330.00~\mathrm{kPa}$
SHSH $16.97~\mathrm{K}$
Fig. 6 shows the Closed-loop state profiles under the MPCs based on RNN, LSTM, LRNN, ICRNN, and ICLRNN for the ORC system. Consistent with the results obtained for the CSTR control case, both ICLRNN-based and ICRNN-based MPC also achieve shorter computation time for the ORC system compared with RNN, LSTM, and LRNN. In particular, the average MPC computation time for the optimal solution with each model is illustrated in Table 9.
Figure 6: Closed-loop state profiles under neural network model-based MPC
Table 9: MPC computation time of neural network models
Model Average computation time (s)
Plain RNN 174.195
Plain LSTM 170.488
LRNN 164.395
ICRNN 133.237
ICL-RNN (ours) 137.672
Specifically, ICLRNN-MPC reduces the average computation time by 20.97%, 19.25%, and 16.26% compared to RNN-MPC, LRNN-MPC, and LSTM-MPC, respectively, while is 3.33% higher than ICRNN-MPC. This superior computational efficiency of ICLRNN and ICRNN further confirms the effectiveness of the introduced convex optimization and highlights their advantage in real-time applicability. Furthermore, since the Lipschitz constraint significantly improves the ICLRNN’s robustness over ICRNN in the ORC modeling study, this careful integration of convexification and gradient restriction serves as a comprehensive development, which enables fast convergence of the optimal solution while preserving robustness, even in the presence of the ORC system with more complexity and stronger nonlinearity.
In general, the proposed ICLRNN outperforms conventional recurrent units in both computational efficiency and robustness through carefully incorporating the convex optimization and the Lipschitz constraint. Its successful application to the modeling and control of engineering scenarios, including the CSTR system and ORC system, demonstrates not only strong noise tolerance but also consistently high computational efficiency during optimal solution. These combined attributes make ICLRNN particularly well-suited for real-time industrial deployment, where both robustness to data uncertainty and rapid solution of optimal control problems are critical for smooth operation.